# Celtibero: Robust Layered Aggregation for Federated Learning

### v1

@@ 1 Introduction
  • We introduce Celtibero, a robust aggregator for Federated Learning (FL) that defends against poisoning attacks, effective in both i.i.d and non-i.i.d scenarios.
@@ 4 Defense Methodology: Celtibero
Celtibero combines detection and filtering with influence reduction to perform robust aggregation. The combination of these two mechanisms enhances the robustness of Celtibero against poisoning attacks. On the one hand, the detection and filtering mechanism aims to eliminate poisoned model updates from the aggregation process. On the other hand, the influence reduction process focus on minimizing the impact of any undetected poisoned updates.
The two mechanisms operate at the layer level. The layered operation is especially relevant for combating attacks that spread through various layers of a model to achieve stealthiness. By accurately identifying and removing a poisoned update at a specific layer, Celtibero can break the poisoning pattern, potentially neutralizing the attack’s impact even if the malicious pattern is not fully detected across all layers. Moreover, the layered approach followed by Celtibero allows for the isolation of layer-specific poisoning patterns, which might be overlooked when considering the entire weight set collectively. Therefore, this approach allows for a more precise and effective defense mechanism with respect to defenses that examine the weights of a specific layer or the weights of all layers in a round.
4.1 Detection, Filterng and Influence Reduction Process
@@
Once gradient vectors have been grouped into two clusters, Celtibero labels the clusters as benign or poisoned based on their size and density. To compute the density of clusters (line 5), Celtibero uses the average pairwise cosine distance between every pair of gradient vectors lying in a cluster. Then, the cluster density is proportionally weighted based on the size of the clusters (line 6). Since poisoned updates will share a common objective, they will tend to form smaller and denser clusters. Therefore, the cluster with the smaller score is determined as posioned by Celtibero, whereas gradient updates lying in the cluster with higher score are considered benign (lines 7-10). Finally, Celtibero performs the aggregation of benign gradient updates using the median values (line 12). The entire process is repeated for each of the layers of the model. The result is a partial aggregation of benign weights $W_{G}^{l^{\prime}}$ at each layer $l$. Finally, all partial layered aggregations are combined to form the global model $W_{G}^{\prime}$.
@@ 5 Experimentation
This section presents the evaluation of Celtibero to show its effectiveness against poisoning attacks in comparison to six state-of-the-art defenses: Median-Krum [25], foolsgold [26], FL-Defender [22], FLAME [16], LFighter [17] and RoseAgg [23]. We conduct all experiments using the Tensorflow deep learning framework [36]. Our code is publicly available at gitlab-borja.
@@ 5.1 Experimental Setup
We consider two typically used metrics in the FL attack literature to evaluate the effectiveness of Celtibero. MTA (Main Task Accuracy) refers to the accuracy of the model on the task it is designed for. It accounts for the proportion of samples that are predicted correctly by the model from all total predictions made by the model. ASR (Attack Success Rate) measures the effectiveness of the poisoning attack under evaluation. For the uLFA, it measures the decay ratio in the MTA accuracy, whereas for the tLFA accounts for the decay on the accuracy of the source class. In the case of backdoor attacks, it indicates the proportion of poisoned samples with the backdoor that are identified by the model as the class intended by the attacker.

### v2

@@ 1 Introduction
  • We introduce Celtibero [^1] Celtibero refers to the ancient tribes that populated the Iberian Peninsula. The Celtiberians are known for their warrior culture and resistance to Roman conquest. They famously fought in the Numantine War, with the siege of Numantia in 133 BC being a notable event where they resisted Rome for an extended period. We named our defense mechanism in honor of their resilience against continuous attacks., a robust aggregator for Federated Learning (FL) that defends against poisoning attacks, effective in both i.i.d and non-i.i.d scenarios.
@@ 4 Defense Methodology: Celtibero
Celtibero combines detection and filtering with influence reduction to perform robust aggregation. The combination of these two mechanisms enhances the robustness of Celtibero against poisoning attacks. The detection and filtering mechanism aims to eliminate poisoned model updates from the aggregation process. This enables robust aggregation since malicious updates typically differ from the majority (benign) updates.
However, some specific attacks may exploit this detection method by crafting updates that resemble benign updates. This aspect makes detection and filtering mechanisms prone to evasion, especially in non-iid environments where benign updates vary from one another. Therefore, putting in risk the global model that is being trained. To address this issue, Celtibero incorporates an influence reduction mechanism that minimizes the impact of undetected poisoned updates. Unlike detection and filtering mechanisms, which average the weights of updates identified as benign, our defense introduces a simple yet robust influence reduction method based on the median. Also, this mechanism provides robustness against updates that slightly modify the weights of models to perform the attack.
Another characteristic of Celtibero is that the two mechanisms operate at the layer level. Contrary to defense approaches that consider the entire set of weights of models, the layered operation of Celtibero is especially relevant for combating attacks that spread through various layers of a model to achieve stealthiness. By accurately identifying and removing a poisoned update at a specific layer, Celtibero can break the poisoning pattern, potentially neutralizing the impact of attacks even if the malicious pattern is not fully detected across all layers. Moreover, the layered approach allows for the isolation of layer-specific poisoning patterns, which might be overlooked when considering the entire weight set collectively. This approach enables a more precise and effective defense compared to mechanisms that only analyze weights from a specific layer or from all layers at once. The next section specifies how all these mechanisms are integrated into Celtibero.
4.1 Layered Detection, Filterng and Influence Reduction Process
@@
Once gradient vectors have been grouped into two clusters, Celtibero scores the clusters according to their size and density to label them as either benign or poisoned. First, Celtibero computes the density of each cluster (line 5) based on the average pairwise cosine distances between every pair of gradient vectors within the cluster. Then, the cluster density is proportionally weighted based on the size of the clusters (line 6). As mentioned, since poisoned updates will share a common attack objective, they will tend to form smaller and denser clusters. Therefore, the cluster with the smaller score is identified as poisoned by Celtibero, whereas gradient updates lying in the cluster with higher score are considered benign (lines 7-10).
Finally, the influence reduction process of Celtibero aggregates the benign gradient updates by taking the median value of each weight (line 12). The result of this operation is a partial aggregation of benign weights $W_{G}^{l^{\prime}}$ at layer $l$ of the model. By repeating the entire process for each of the individual layers of the model, Celtibero performs layered aggregation. Finally, the partial aggregations for all layers are combined to form the global model $W_{G}^{\prime}$ update.
@@ 5 Experimentation
This section presents the evaluation of Celtibero to show its effectiveness against poisoning attacks in comparison to six state-of-the-art defenses: Median-Krum [25], foolsgold [26], FL-Defender [22], FLAME [16], LFighter [17] and RoseAgg [23]. We conduct all experiments using the Tensorflow deep learning framework [36]. Our code is publicly available at gitlab-borja [^2] We will publish the link upon paper acceptance.
@@ 5.1 Experimental Setup
We consider two typically used metrics in the FL attack literature to evaluate the effectiveness of Celtibero. MTA (Main Task Accuracy) refers to the accuracy of the model on the task it is designed for. It accounts for the proportion of samples that are predicted correctly by the model from all total predictions made by the model. ASR (Attack Success Rate) measures the effectiveness of the poisoning attack under evaluation. Since label flipping attacks seek to harm the detection capacity of the model, for the uLFA, the ASR measures the decay ratio in the MTA accuracy with respect to the no-attack model, whereas for the tLFA, the ASR accounts for the decay on the accuracy of the source class with respect to the corresponding value of the no-attack model. In the case of backdoor attacks, the ASR indicates the proportion of poisoned samples with the backdoor that are identified by the model as the class intended by the attacker.